I have a chapter 5 test review and have gone to my math teacher for help today but it still did not help. I do not understand what a function is. He says that f isn't important. But that just confused me even more. The question is "Write the function that has the given values in slope-intercept form." The numbers are f(0)=3, f(-2)=-7. If you could help me this would be great. I am in 9th grade and take algebra 1. Thank you for your time!
Answer by richard1234(7193)

The letter f defines the function itself, so it is indeed important.

A function, in general terms, is defined as an operation that "maps" a set of values (called the "domain") to another set of values (called the "range"). Usually, the domain and range will be the set of real numbers (unless there are certain restrictions). Eventually you will see functions where the domain and/or range require multiple numbers or variables.

For example, suppose f(x) = 2x + 1. Then f(4) = 2(4) + 1 = 9 and f(-1.5) = 2(-1.5) + 1 = -2.

To solve the question you posted, you will have to know what "slope" is. Slope is essentially how "steep" the function is if it were to be graphed, and it is usually denoted by the letter m. The slope between two points (x1,y1), (x2,y2) is



So just plug in the points (0,3) and (-2,7) into the formula. You already know the y-intercept is (0,3) so all you need is the slope.